# Locally recoverable algebro-geometric codes with multiple recovery sets from projective bundles

### v1

@@ 1.1. Algebro-geometric context
In 1972, Justesen [Justesen72] pioneered the systematic study of families of codes with good asymptotic behavior; his own codes extended Reed–Solomon codes but did not use algebro-geometric techniques. Shortly after Goppa introduced algebro-geometric methods, Katsman, Tsfasman and Vlăduţ constructed algebro-geometric families of codes with good asymptotic properties [KatsmanTsfasmanVluadact84]. Further constructions appeared over time, e.g. [vLintSpringer87], including recent work using algebraic surfaces [CouvreurLebacquePerret21]. It is natural to wonder if it’s possible to construct algebro-geometric families of asymptotically good codes that incorporate locality and availability; to date, there are surprisingly few constructions of such families [BargTamoVluadact17, LiLiuMaXing24].
@@ 1.2. Results
Let $b\geq 3$ and $r=1,2,3$. The code $\mathcal{C}$ is optimal and locally recoverable with locality $r$. Its parameters $[n,k,d]_{q}$ are
@@
Our second main result is that when $q$ is large enough, for most choices of ${\mathcal{P}}$, the code ${\mathcal{C}}$ is optimal and locally recoverable.
@@
Let $b\geq 3$ and $r\geq 4$. There exists an integer $q_{0}=q_{0}(r,b)$ such that if $q\geq q_{0}$, for most choices of points ${\mathcal{P}}$ there are no code words in ${\mathcal{C}}$ of weight $\leq r+2$. That is, the minimum distance of ${\mathcal{C}}$ is $d\geq r+3$. Consequently, for most choices of points ${\mathcal{P}}$ the code ${\mathcal{C}}$ is optimal and locally recoverable with locality $r$. Moreover, as $q\to\infty$, the code ${\mathcal{C}}$ is optimal with probability $1$.
@@
The cardinality of a set $S$ will be denoted by #$\#S$. A code will be denoted by a calligraphic letter ${\mathcal{C}}$, and its dual code will be denoted by ${\mathcal{C}}^{\bot}$. All the codes we handle are linear evaluation codes ${\mathcal{C}}\coloneqq\mathrm{im}\operatorname{ev}_{{\mathcal{P}}}$ with ${\mathcal{P}}$ a finite set of points. A code word of ${\mathcal{C}}$ will be denoted by a bold lowercase letter $\mathbf{c}$. Scalars in ${\mathbb{F}}_{q}$ will be denoted by a lowercase $x$, and elements of ${\mathbb{F}}_{q}^{n}$ for $n>1$ will be denoted by a bold lowercase $\mathbf{x}$. The $n$-dimensional affine space over ${\mathbb{F}}_{q}$ is denoted by ${\mathbb{A}}^{n}$, and its Cartesian coordinates are denoted by ${\mathbb{A}}_{x_{1},\dots,x_{n}}^{n}$. The $n$-dimensional projective space over ${\mathbb{F}}_{q}$ is denoted by ${\mathbb{P}}^{n}$, its homogeneous coordinates are denoted by ${\mathbb{P}}_{\mathbf{x}}^{n}$ where $\mathbf{x}=[x_{0},\dots,x_{n}]$. When we write ${\mathbb{P}}_{x_{0},\dots,x_{n-1}}^{n}$, the coordinates $x_{0},\dots,x_{n-1}$ mean affine coordinates in the chart $D_{+}(x_{n})$ where the last homogeneous coordinate $x_{n}$ is nonzero. Given a projective variety $X$ over ${\mathbb{F}}_{q}$, we denote by $X({\mathbb{F}}_{q})$ the ${\mathbb{F}}_{q}$-rational points of $X$.
@@ 3.1. Set-up
$A_{2}:=\{(6,5),(6,6),(6,7),(6,8)\}\quad\text{and}\quad A_{3}:=\{(23,20),(23,21),(23,22),(23,23)\}$
@@ 3.2. Relations to existing codes
A code ${\mathcal{C}}$ as in Section 3.1 has important similarities and differences with the codes considered in [TamoBarg14, SalgadoVarillyAlvaradoVoloch21]. In [SalgadoVarillyAlvaradoVoloch21]*III.B, the authors consider algebro-geometric codes arising from a triple of data $(X,{\mathcal{P}},V[N])$ where $X={\mathbb{A}}^{1}_{x}\times{\mathbb{A}}^{1}_{y}$, like our codes, and where
@@
for a nonnegative integer $N$. Our codes consider mostly the case where $N=b-2$, though we look at smaller values of $N$ in Section 3.7. The key difference between their codes and ours is the choice of points ${\mathcal{P}}$ for code evaluation. We all consider a set of points ${\mathcal{P}}$ partitioned into $b$ batches $A_{1},\dots,A_{b}$ consisting of $r+1$ distinct points. However, in [SalgadoVarillyAlvaradoVoloch21]*III.B each batch is additionally constrained to satisfy an extra algebraic relation: If $A_{i}=\{(x_{i},y_{i,j})\}_{j=1}^{r+1}$, then there exists $g(x)\in{\mathbb{F}}_{q}[x]$ a polynomial of degree $r+1$ such that $y_{i,j}=g(x_{i})$ for all $1\leq i\leq b$ and $11\leq j\leq r+$ 1. Tamo and Barg’s construction in [TamoBarg14] uses the polynomial $g(x)=x^{r+1}$, and more generally discusses the concept of good polynomials $g(x)$. The polynomial $g(x)$ affords better control of the minimum distance of the resulting codes. Without this kind of control, we can only prove that the codes constructed in Section 3.1 are optimal when $b\geq 3$ and $r=1$, $2$ or $3$; see Theorem 3.13. However, when $r\geq 4$, our codes in Section 3.1 are provably optimal with probability approaching $1$ for uniformly random choices of points ${\mathcal{P}}$ as $q\to\infty$; see Theorem 3.15. Although our codes for $r\geq 4$ are not always optimal (see Section 4), our investigation suggests that the use of good polynomials, like $g(x)=x^{r+1}$, imposes a serious constraint on the universe’s supply of optimal LRCs.
@@ 3.6. A detour through matroids
so that $MD\mathbf{v}^{\prime}=M\mathbf{v}=0$. Thus $\ker MB=\ker M^{\prime}$, so the row spaces of $MD$ and $M^{\prime}$ coincide, so by doing elementary row operations to $MD$ we can reach $M^{\prime}$. The matrix $A$ encoding those elementary row operations satisfies $AMD=M^{\prime}$ as claimed. ∎
@@
To show that the identity map on groundsets induces an isomorphism of matroids between $\mathsf{M}_{{\mathcal{C}}}$ and $\mathsf{M}_{{\mathcal{C}}^{\prime}}$ it suffices to show that we can get from $G$ to $G^{\prime}$ via a sequence of row operations and column scaling. That is, we want to find an invertible $k\times k$ matrix $T$ and a diagonal $n\times n$ matrix $R$ satisfying $TGR=G^{\prime}$. Let $S=\{x_{1},\dots,x_{b}\}$ and $S^{\prime}=\{x^{\prime}_{1},\dots,x^{\prime}_{b}\}$, and let $A$ and $D=\mathrm{diag}(D_{1},\dots,D_{b})$ be as in Lemma 3.21. Consider
@@ 3.7. Constructing codes with larger minimum distance
$V_{z}:=\left\{\sum_{\ell=0}^{r-1}a_{\ell}(x)y^{\ell}:a_{\ell}(x)\in{\mathbb{F}}_{q}[x],\deg a_{\ell}(x)\leq b-2-z\right\}\subset{\mathbb{F}}_{q}[x,y].$
@@
${\mathcal{C}}_{z}={\mathcal{C}}_{z}(X,{\mathcal{P}},V_{z}):=\operatorname{im}\left(\operatorname{ev}_{\mathcal{P}}\colon V_{z}\to{\mathbb{F}}_{q}^{n}\right),$
@@ 5. Codes from projective space bundles
Let $f(x,\mathbf{y})$ be a nonzero polynomial in $V$. Then $f(x_{i},\mathbf{y}_{i,j})=0$ for all $\mathbf{y}_{i,j}\in A_{i}$ for $s\leq\alpha$ values of $x_{1},\dots,x_{b}$, by Lemma 5.4. If $i\in\{1,\dots,b\}$ is such that $f(x_{i},\mathbf{y}_{i,j})\neq 0$ for at least one $\mathbf{y}_{i,j}\in A_{i}$, then there are $\leq r-1$ points $\mathbf{y}_{i,j_{1}},\dots,\mathbf{y}_{i,j_{r-1}}\in A_{i}$ such that $f(x_{i},\mathbf{y}_{i,j_{1}})=\cdots=f(x_{i},\mathbf{y}_{i,j_{r-1}})=0$ because the points in $A_{i}$ are in general position. Consequently the number of zeros of $f(x,\mathbf{y})$ along each of the $b-s$ non zero-fibers is $\geq(tr+1)-(r-1)$, and thus the weight $\omega$ of the code word $\mathbf{c}:=\operatorname{ev}_{{\mathcal{P}}}(f(x,\mathbf{y}))$ is bounded by
@@ 5.1. Gilbert–Varshamov bounds
Given an asymptotically good family of codes $\mathscr{C}$ like the one produced above, it is natural to ask if the limiting relative parameters in the proof of Theorem 5.2 lie near a Gilbert–Varshamov-type bound for LRCs with availability. Barg, Tamo, and Frolov obtained a bound for LRCs with availability in [TamoBargFrolov16]*Theorem B. It is difficult to derive an asymptotic bound from their formulas, although they do this successfully in the case of availability $t=2$ [TamoBargFrolov16]*Figure 1. Their asymptotic bounds, as $n\to\infty$, hold the availability $r$ fixed. Unfortunately, for the codes in our family, $n$ and $r$ are inextricably linked; if $n\to\infty$ then $r\to\infty$ as well. Therefore, our family of codes $\mathscr{C}$ is not amenable to an asymptotic GV-bound analysis. Exploring whether an alternative construction could be analyzed from this perspective would be worthwhile.

### v2

# Locally recoverable algebro-geometric codes from projective bundles

@@ 1.1. Algebro-geometric context
In 1972, Justesen [Justesen72] pioneered the systematic study of families of codes with good asymptotic behavior; his own codes extended Reed–Solomon codes but did not use algebro-geometric techniques. Shortly after Goppa introduced algebro-geometric methods, Katsman, Tsfasman, and Vlăduţ constructed algebro-geometric families of codes with good asymptotic properties [KatsmanTsfasmanVluadact84]. Further constructions appeared over time, e.g. [vLintSpringer87], including recent work using algebraic surfaces [CouvreurLebacquePerret21]. It is natural to wonder if it’s possible to construct algebro-geometric families of asymptotically good codes that incorporate locality and availability; to date, there are surprisingly few constructions of such families [BargTamoVluadact17, LiLiuMaXing24].
@@ 1.2. Results
Let $r=1$, $2$, or $3$ and let $b$ satisfy $3\leq b\leq\frac{q}{r+1}$. The code $\mathcal{C}$ is optimal and locally recoverable with locality $r$. Its parameters $[n,k,d]_{q}$ are
@@
This result is best possible in the following sense: For any $q$, if $r\geq 4$, we can construct nonoptimal codes for all $b$ satisfying $3\leq b\leq\frac{q}{r+1}$. We do this concretely as a proof of concept in Section 4 for $r=b=4$ and $q=37$. Knowing that not all of our codes are optimal, we naturally ask what proportion of these codes are optimal. Our second main result answers this question: When $q$ is large enough, for most choices of ${\mathcal{P}}$, the code ${\mathcal{C}}$ is optimal and locally recoverable. In other words, almost all of the codes we construct are optimal.
@@
Let $r\geq 4$ and let $b$ satisfy $3\leq b\leq\frac{q}{r+1}$. There exists an integer $q_{0}=q_{0}(r,b)$ such that if $q\geq q_{0}$, for most choices of points ${\mathcal{P}}$ there are no code words in ${\mathcal{C}}$ of weight $\leq r+2$. That is, the minimum distance of ${\mathcal{C}}$ is $d\geq r+3$. Consequently, for most choices of points ${\mathcal{P}}$ the code ${\mathcal{C}}$ is optimal and locally recoverable with locality $r$. Moreover, as $q\to\infty$, the code ${\mathcal{C}}$ is optimal with probability $1$.
@@
It is worth pointing out that the codes we construct to prove Theorem 1.2 generalize the optimal LRCs of Tamo and Barg in [TamoBarg14], while avoiding the use of good polynomials. A noteworthy difference between our codes and theirs is that we do not require evaluation points to be carefully chosen; a random choice of points yields an optimal LRC with probability $\approx 1$, as long as $q$ is large (see Remark 3.16 for more details).
@@
The cardinality of a set $S$ will be denoted by #$\#S$. A code will be denoted by a calligraphic letter ${\mathcal{C}}$, and its dual code will be denoted by ${\mathcal{C}}^{\bot}$. All the codes we handle are linear evaluation codes ${\mathcal{C}}\coloneqq\mathrm{im}\operatorname{ev}_{{\mathcal{P}}}$ with ${\mathcal{P}}$ a finite set of points. A code word of ${\mathcal{C}}$ will be denoted by a bold lowercase letter $\mathbf{c}$. Scalars in ${\mathbb{F}}_{q}$ will be denoted by a lowercase $x$, and elements of ${\mathbb{F}}_{q}^{n}$ for $n>1$ will be denoted by a bold lowercase $\mathbf{x}$. The $n$-dimensional affine space over ${\mathbb{F}}_{q}$ is denoted by ${\mathbb{A}}^{n}$, and its Cartesian coordinates are denoted by ${\mathbb{A}}_{x_{1},\dots,x_{n}}^{n}$. The $n$-dimensional projective space over ${\mathbb{F}}_{q}$ is denoted by ${\mathbb{P}}^{n}$, its homogeneous coordinates are denoted by ${\mathbb{P}}_{\mathbf{x}}^{n}$ where $\mathbf{x}=[x_{0},\dots,x_{n}]$. When we write ${\mathbb{P}}_{x_{0},\dots,x_{n-1}}^{n}$, the coordinates $x_{0},\dots,x_{n-1}$ are the affine coordinates in the chart $D_{+}(x_{n})$ where the last homogeneous coordinate $x_{n}$ is nonzero. Given a projective variety $X$ over ${\mathbb{F}}_{q}$, we denote by $X({\mathbb{F}}_{q})$ the ${\mathbb{F}}_{q}$-rational points of $X$.
@@ 3.1. Set-up
$A_{2}\coloneqq\{(6,5),(6,6),(6,7),(6,8)\}\quad\text{and}\quad A_{3}\coloneqq\{(23,20),(23,21),(23,22),(23,23)\}$
@@ 3.2. Relations to existing codes
A code ${\mathcal{C}}$ as in Section 3.1 has important similarities and differences with the codes considered in [TamoBarg14, SalgadoVarillyAlvaradoVoloch21]. In [SalgadoVarillyAlvaradoVoloch21]*Section III.B, the authors consider algebro-geometric codes arising from a triple of data $(X,{\mathcal{P}},V[N])$ where $X={\mathbb{A}}^{1}_{x}\times{\mathbb{A}}^{1}_{y}$, like our codes, and where
@@
for a nonnegative integer $N$. Our codes consider mostly the case where $N=b-2$, though we look at smaller values of $N$ in Section 3.7. The key difference between their codes and ours is the choice of points ${\mathcal{P}}$ for code evaluation. We all consider a set of points ${\mathcal{P}}$ partitioned into $b$ batches $A_{1},\dots,A_{b}$ consisting of $r+1$ distinct points. However, in [SalgadoVarillyAlvaradoVoloch21]*Section III.B each batch is additionally constrained to satisfy an extra algebraic relation: If $A_{i}=\{(x_{i},y_{i,j})\}_{j=1}^{r+1}$, then there exists $g(x)\in{\mathbb{F}}_{q}[x]$ a polynomial of degree $r+1$ such that $y_{i,j}=g(x_{i})$ for all $1\leq i\leq b$ and $11\leq j\leq r+$ 1. Tamo and Barg’s construction in [TamoBarg14] uses the polynomial $g(x)=x^{r+1}$, and more generally discusses the concept of good polynomials $g(x)$. The polynomial $g(x)$ affords better control of the minimum distance of the resulting codes. Without this kind of control, we can only prove that the codes constructed in Section 3.1 are optimal when $b\geq 3$ and $r=1$, $2$ or $3$; see Theorem 3.13. However, when $r\geq 4$, our codes in Section 3.1 are provably optimal with probability approaching $1$ for uniformly random choices of points ${\mathcal{P}}$ as $q\to\infty$; see Theorem 3.15. Although our codes for $r\geq 4$ are not always optimal (see Section 4), our investigation suggests that the use of good polynomials, like $g(x)=x^{r+1}$, imposes a serious constraint on the universe’s supply of optimal LRCs.
Our codes also share superficial similarities with those of Munuera and Tenório [MunueraTenorio18]*Section 2.2, which is not surprising, as their codes generalize those in [TamoBarg14, BargTamoVluadact17]. However, our codes are neither a special case of the Munuera–Tenório construction, nor is there a clear common refinement of both constructions. To mimic our codes in the notation of [MunueraTenorio18], one would need to take $m=2$, $t=1$, $\phi_{1}=x_{1}\eqqcolon x$, $\phi_{2}=x_{2}\eqqcolon y$, #$\#{\mathcal{S}}=b$, and $V_{i}=\{a(x)\in{\mathbb{F}}_{q}[x]:\deg a(x)\leq b-2\}$ for all $10\leq i\leq r-$ 1. This would force $r=q-1$ in their set-up (a restriction we do not impose), and the set of evaluation points ${\mathcal{P}}\subset{\mathbb{A}}^{2}_{x,y}({\mathbb{F}}_{q})$ would contain $b(r+1)=bq$ points in $b$ batches with $r+1=q$ overlapping $y$ coordinates. In our construction, it is essential that all the $b(r+1)$ points have distinct $y$ coordinates; see Remarks 3.1(3) and Figure 1.
@@ 3.6. A detour through matroids
so that $MD\mathbf{v}^{\prime}=M\mathbf{v}=0$. Thus $\ker MD=\ker M^{\prime}$, so the row spaces of $MD$ and $M^{\prime}$ coincide, so by doing elementary row operations to $MD$ we can reach $M^{\prime}$. The matrix $A$ encoding those elementary row operations satisfies $AMD=M^{\prime}$ as claimed. ∎
@@
To show that the identity map on groundsets induces an isomorphism of matroids between $\mathsf{M}_{{\mathcal{C}}}$ and $\mathsf{M}_{{\mathcal{C}}^{\prime}}$ it suffices to show that we can get from $G$ to $G^{\prime}$ via a sequence of row operations and column scaling. That is, we want to find an invertible $k\times k$ matrix $T$ and a diagonal $n\times n$ matrix $R$ satisfying $TGR=G^{\prime}$. Let $S=\{x_{1},\dots,x_{b}\}$ and $S^{\prime}=\{x^{\prime}_{1},\dots,x^{\prime}_{b}\}$, and let $A$ and $D=\mathrm{diag}[D_{1},\dots,D_{b}]$ be as in Lemma 3.21. Consider
@@ 3.7. Constructing codes with larger minimum distance
$V_{z}\coloneqq\left\{\sum_{\ell=0}^{r-1}a_{\ell}(x)y^{\ell}:a_{\ell}(x)\in{\mathbb{F}}_{q}[x],\deg a_{\ell}(x)\leq b-2-z\right\}\subset{\mathbb{F}}_{q}[x,y].$
@@
${\mathcal{C}}_{z}={\mathcal{C}}_{z}(X,{\mathcal{P}},V_{z})\coloneqq\operatorname{im}\left(\operatorname{ev}_{\mathcal{P}}\colon V_{z}\to{\mathbb{F}}_{q}^{n}\right),$
@@ 5. Codes from projective space bundles
Let $f(x,\mathbf{y})$ be a nonzero polynomial in $V$. Then $f(x_{i},\mathbf{y}_{i,j})=0$ for all $\mathbf{y}_{i,j}\in A_{i}$ for $s\leq\alpha$ values of $x_{1},\dots,x_{b}$, by Lemma 5.4. If $i\in\{1,\dots,b\}$ is such that $f(x_{i},\mathbf{y}_{i,j})\neq 0$ for at least one $\mathbf{y}_{i,j}\in A_{i}$, then there are $\leq r-1$ points $\mathbf{y}_{i,j_{1}},\dots,\mathbf{y}_{i,j_{r-1}}\in A_{i}$ such that $f(x_{i},\mathbf{y}_{i,j_{1}})=\cdots=f(x_{i},\mathbf{y}_{i,j_{r-1}})=0$ because the points in $A_{i}$ are in general position. Consequently the number of zeros of $f(x,\mathbf{y})$ along each of the $b-s$ non zero-fibers is $\geq(tr+1)-(r-1)$, and thus the weight $\omega$ of the code word $\mathbf{c}\coloneqq\operatorname{ev}_{{\mathcal{P}}}(f(x,\mathbf{y}))$ is bounded by
@@ 5.1. Gilbert–Varshamov bounds
Given an asymptotically good family of codes $\mathscr{C}$ like the one produced above, it is natural to ask if the limiting relative parameters in the proof of Theorem 5.2 lie near a Gilbert–Varshamov-type bound for LRCs with availability. Barg, Tamo, and Frolov obtained a bound for LRCs with availability in [TamoBargFrolov16]*Theorem B. It is difficult to derive an asymptotic bound from their formulas, although they do this successfully in the case of availability $t=2$ [TamoBargFrolov16]*Figure 1. Their asymptotic bounds, as $n\to\infty$, hold the locality $r$ fixed. Unfortunately, for the codes in our family, $n$ and $r$ are inextricably linked; if $n\to\infty$ then $r\to\infty$ as well. Therefore, our family of codes $\mathscr{C}$ is not amenable to an asymptotic GV-bound analysis. Exploring whether an alternative construction could be analyzed from this perspective would be worthwhile.